2 pt Question 12 A project is being funded by a 6% tax-free 20 year bond. Inflation is expected to average 2% annually over this time period. The resulting Rate of Return for the bond should be what? 5.88% 1.96% 3.65% 3.92% @ 4.00% Previous Next
Answer #1

Answer 12

The correct answer is (d) 3.92%

Formula:

Real Rate of Return is given by:

r = (1 + i)/(1 + e) - 1

where i = Nominal rate of return = 6% = 0.06

e = inflation rate = 2% = 0.02

r = real rate of return

Hence, r = (1 + 0.06)(1 + 0.02) - 1 = 0.0392 ~ 3.92%

Hence real rate of return = 3.92%

Hence, the correct answer is (d) 3.92%